2x[sup]3[/sup] + 3x[sup]2[/sup] - 8x -12 = 0& (given that 2 is a root)
1
Expert's answer
2011-03-16T10:02:23-0400
x1 = 2
(2x3 + 3x2 - 8x - 12)/(x-2) = 2x2 + 7x +6
2x2 + 7x +6 = 0
D = 49 - 48 = 1
x2 = -1.5
x3 = -2